INTRACAVITY LASER SPECTROSCOPY OF NiCl: SYSTEM F

\underline{JAMES OBRIEN}, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499; and LEAH OBRIEN, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The red electronic transition of NiCl occurring in the region of 14430 cm^{−1}, also known as System F, has been recorded with rotational resolution by intracavity laser absorption spectroscopy. The gas phase NiCl molecules were produced in an electric discharge using a nickel cathode and a trace amount of carbon tetrachloride with approx. 1 torr argon. Results of the analysis will be presented.